Two rectangles are similar. The ratio of the lengths of their corresponding sides is 3:4. Find the ratios of the areas of the two rectangles.
Answer by SmurfDude(4)

A RATIO is the quotient of two numbers or two quantities.
The RATIO of 3:4 can also be written as:
3/4 or 3:4
A PROPORTION is a mathematical statement that two ratios are equal.
In the proportion a/b = c/d, the products ad and bc are called cross products.
If a/b = c/d then ad=bc
in your question a ratio of 3:4 could be exampled as:
one rectangle could be measured as 3" two sides and 4" on the other two
giving you your original 3:4 ratio
another example would be a rectangle measuring at 12 inches by 16 inches:
Here is your RATIO PROPORTION:
3/4 = 12/16
You get the similar ratios by multiplying the numerator of the first ratio by the denominator of the second ratio which would be 3 * 16 = 48
and 4 * 12 = 48
There you have an example of a 3:4 ratio